OpenStudy (anonymous):

I will give away a award to whoever help me on this question. Solve: x/10-5=2/5

OpenStudy (smileyxl3):

You want to find x?

OpenStudy (anonymous):

yes

OpenStudy (smileyxl3):

ok, \[\frac{ x }{10 }=\frac{ 27 }{5 }\]

OpenStudy (smileyxl3):

Cross-multiply it

OpenStudy (anonymous):

Some of my answers are 7,9,25,54

OpenStudy (smileyxl3):

(10*27)/5=54